OpenStudy (mineymouse):

Please help!! will give medals!! Mr. Hewitt is hoping that Mrs. Reed will buy him a geometrically interesting piece of land on which he will build the most fabulous castle the world has ever seen. Mrs. Reed has said that she will only buy the land if it costs less than $4.00 per square foot. The plot shown below is listed for $453,175. Is Mr. Hewitt going to be able to get this land?

OpenStudy (ericamarie7):

Show the diagram?

OpenStudy (ericamarie7):

The answer is Yes.

OpenStudy (mineymouse):

can you teach me how you arrived to that though?

OpenStudy (anonymous):

\[\frac{453175}{\frac{1}{2} (390-225) (187.5\, +397.5)+\frac{225 (225+397.5)}{2}}=3.83093 \]